OpenStudy (anonymous):

What is the conversion factor between ft2 and yd2?

OpenStudy (jchick):

Square the NOT values. Your answer will appear.

OpenStudy (jchick):

9 ft^2 = 1yd^2 b) (3.281)^2 ft^2 = m^2 10.765 ft^2 = m^2